OpenStudy (anonymous):

Solve the equations. -(7y+1)-(-6y-3)=-4

OpenStudy (saifoo.khan):

\[-7y -1 +6y+3 = -4\]

OpenStudy (saifoo.khan):

Solve.

OpenStudy (saifoo.khan):

-y + 2 = -4 -y = -6 y = 6